OpenStudy (anonymous):

what is the equation of this (-2/-3), (4/5)?

Nnesha (nnesha):

is it (-2,-3)(4,5) ??

OpenStudy (anonymous):

@Nnesha no i have options to choose from if you would like to see them

Nnesha (nnesha):

yea i would like to se

OpenStudy (anonymous):

A. y=4/3x-1/3 B y= -4/3x-1/3 C y=3/4x-1/4 D y=-4/3x-1/4

Nnesha (nnesha):

is it look like this ?\[(\frac{ -2 }{ -3 }) ,(\frac{ 4 }{ 5})\]?

OpenStudy (popitree):

the slope m is (y2-y1)/(x2-x1). And equation is, y=mx+c. Solve for c with putting any point in the equation. or y-y1 = m(x-x1)

Nnesha (nnesha):

yeah but there supposed to be two points (x,y) (x,y)